# EE-MLLM: A Data-Efficient and Compute-Efficient Multimodal Large Language Model

### v1

@@ Introduction
We conduct our evaluations using the VLMEvalKit (Duan et al. 2024), and the results of other state-of-the-art models are obtained from the same source.
@@
General Benchmarks. 1) MMBench-EN (Liu et al. 2023c) is a comprehensive multimodal benchmark specifically designed to assess the performance of MLLMs. It comprises over 3,000 multiple-choice questions spanning 20 ability categories. We evaluate EE-MLLM on the MMBench-EN-V1.1. 2) MME (Fu et al. 2023) evaluates advanced MLLMs in terms of perception and cognition, encompassing a total of 14 subtasks. To minimize the impact of prompt engineering on MLLMs, MME’s instructions are designed to elicit simple binary responses, such as ”please answer yes or no.”. We report the results of the perception split of MME. 3) ScienceQA (Lu et al. 2022)is derived from elementary and high school science curricula. The questions in ScienceQA cover three subjects: natural science, language science, and social science. 4) HallusionBench (Guan et al. 2023) is designed for evaluating image-context reasoning, comprising 346 images paired with 1,129 questions crafted by human experts. HallusionBench takes into account both language hallucinations and visual illusions across a diverse range of topics. 5) MMMU (Yue et al. 2023) collects 11.5K multimodal questions from college exams, quizzes, and textbooks, covering six core disciplines, spanning 30 subjects and 183 subfields, including 30 heterogeneous image types. 6) CCBench (Liu et al. 2023c), developed by the MMBench team, is specifically designed to evaluate MLLMs in the domain of Chinese Culture. 7) SeedBench (Li et al. 2023) encompasses 19K multiple choice questions, covering 12 evaluation dimensions, including both image and video. We only use questions with images for evaluation. 8) BLINK (Fu et al. 2024) contains 14 visual perception tasks which pose significant challenges for current multimodal LLMs.
Model LLM AI2D OCRBench TextVQA ChartQA DocVQA Seed2 Plus
InstructBLIP Vicuna-7B 40.6 27.6 33.6 10.9 - 29.5
MiniGPT-4-v1 Vicuna-7B 28.4 17.2 - - - 15.2
MiniGPT-4-v2 Llama2-7B 30.5 3.1 - - - 23.3
Idefics-instruct Llama-7B 42.2 25.2 - - - 35.4
OpenFlamingo v2 MPT-7B 31.7 14.9 16.3 - - 28.7
Qwen-VL Qwen-7B 57.7 12.7 63.1 59.0 - 40.1
Qwen-VL-Chat Qwen-7B 63.0 48.8 60.7 49.8 - 40.6
ShareGPT4V Vicuna-7B 58.0 37.1 51.1 21.3 - 46.1
LLaVA-v1.5 Vicuna-7B 55.5 31.8 45.5 17.8 - 41.3
LLaVA-v1.6 Vicuna-7B 67.0 53.2 64.4 55.4 - 51.6
EE-MLLM Vicuna-7B 65.7 57.4 69.0 68.4 67.4 53.5
Table 3: Comparsion with state-of-the-art methods on fine-grained benchmarks, including AI2D, OCRBench, TextVQA, ChartQA, DocVQA, InfoVQA, and SeedBench-2 Plus.
Model MMB-VAL SeedBench ScienceQA BLINK AI2D OCRBench TextVQA ChartQA Avg.
$336\times 336$
LLaVA-v1.5 67.3 65.1 67.0 40.7 56.4 24.4 46.1 17.9 48.1
EE-MLLM 66.8 64.1 66.7 38.4 56.2 23.6 43.3 17.3 47.1
$672\times 672$
LLaVA-v1.5 70.1 68.0 69.4 40.1 58.3 39.8 61.7 27.9 54.4
EE-MLLM 68.2 67.3 69.1 40.3 57.5 37.0 60.1 25.1 53.1
Table 4: Comparisons with LLaVA under the same settings, including LLM, vision encoder and training data.
Fine-grained Benchmarks. 1) AI2D (Kembhavi et al. 2016) emphasizes diagram interpretation and reasoning, comprising 5,000 diagrams and 15,000 questions and answers. 2) OCRBench (Liu et al. 2023d) aims to facilitate the assessment of MLLM OCR capabilities, including 29 datasets. 3) TextVQA (Singh et al. 2019) consists of 45,336 questions and 28,408 images that necessitate reasoning about text for answering. We use the validation set, containing 5,000 images, for evaluation. 4) ChartQA (Masry et al. 2022) is a large-scale benchmark featuring 20,882 charts, with questions focusing on logical and visual reasoning. 5) DocVQA (Mathew, Karatzas, and Jawahar 2021) concentrates on document image understanding, encompassing 50,000 questions and over 12,000 images. We evaluate using the validation set, which includes 5,349 questions and 1,286 images. 6) Seed2 Plus (Li et al. 2024a) specifically designed for text-rich visual comprehension evaluation of MLLMs, includes 2.3K multiple-choice questions covering charts, maps, and webs.
Figure 3: The visualization results of examples sampled from BLINK and RealWorldQA.
Comparisons with State-of-the-art Models
General Benchmarks. In Tab. 2, we compare EE-MLLM with various state-of-the-art MLLMs (Liu et al. 2024; Bai et al. 2023; Ma et al. 2024), across eight general benchmarks. These benchmarks evaluate the comprehensive ability of MLLMs, including understanding and perception, as well as the severity of illusion. These factors collectively reflect the generalization ability and applicability of MLLMs in real-world scenarios. EE-MLLM achieves comparable performance with state-of-the-art MLLMs in general benchmarks. Specifically, EE-MLLM achieves a score of 70.4 in MMBench and 1528.1 in MME, scores that are notably higer compared to LLaVA-v1.6, which also support high-resolution images input. This suggests that EE-MLLM possesses comprehensive perception and reasoning capabilities. What’s more, EE-MLLM also achieves promising performance on CCBench and SeedBench.
@@
Figure 4: Illustration of Inference Speed Ratio between EE-MLLM and LLaVA for input resolutions of $980\times 980$, considering numbers of generated tokens ranging from 2 to 256, for single and double image inputs.
Fine-grained Benchmarks. In Tab. 3, we perform evaluations on seven fine-grained benchmarks. These benchmarks necessitate impressive visual perception capability of MLLMs, as they require the exploration of fine-grained information within images to answer the questions. For traditional VQA benchmarks, such as TextVQA, ChartQA, and DocVQA, MLLM achieves a very impressive performance, 4.6 higher on TextVQA and 13.0 higher on ChartQA compared to LLaVA-v1.6, which also supports high resolution. In the OCRBench, specifically designed to assess the OCR capabilities of MLLM, EE-MLLM outperforms by 4.2 compared to LLaVA-v1.6. These results demonstrate that despite EE-MLLM significantly reducing the computational overhead associated with visual tokens, it effectively maintains the model’s fine-grained capabilities.
Comparisons of Inference Speed
@@
Despite the significant reduction in FLOPs demonstrated by EE-MLLM, it is crucial to consider real-world scenarios where advanced techniques such as KVCache and Batch Inference are typically employed during deployment. Hence, conducting comparisons of inference speed with these techniques becomes necessary. We conduct the comparisons of inference speed on a single NVIDIA H800. The resolution of the input image is set to $980\times 980$, and the number of generated tokens varies from 2 to 256. We illustrate the speed ratio between EE-MLLM and LLaVA in Fig. 4. Our findings show that when generating 8 tokens, EE-MLLM’s inference speed is three times that of LLaVA. However, as the number of generated tokens increases, the speed ratio decreases. When generating 64 tokens, EE-MLLM’s inference speed is 1.6 times that of LLaVA. The reason for this observation is that our EE-MLLM primarily reduces the computational cost during the prefilling stage, which calculates the KV cache of visual tokens. The generation of the first token is faster than self-attention-based methods like LLaVA. However, the advantage in inference speed diminishes after the first token. Specifically, with two input images, the inference speed of EE-MLLM is nearly four times that of LLaVA. This clearly indicates that EE-MLLM is significantly more efficient in multi-image input scenarios, including interleaved image-text dialogues and multimodal in-context learning.
Stage EE-MLLM LLaVA
Pre-training 32G 75G
Fine-tuning 66G 69G
Table 5: Memory usage of GPU during pre-training and fine-tuning stages.
GPU Memory Overhead of EE-MLLM
We assess the GPU memory overhead of EE-MLLM during both the pre-training and fine-tuning stages. Similarly, we compare EE-MLLM with LLaVA at a resolution of $336\times 336$. We perform experiments on 8$\times$×H800, with a global batch size of 256 for pre-training, 128 for fine-tuning. The comparison of memory usage is detailed in Tab 5. During the pre-training stage, EE-MLLM exhibits significant lower memory usage, at 32G, while the memory usage of LLaVA is 75G. During the fine-tuning stage, as the primary memory usage is focused on the trainable LLM, the memory usage advantage becomes less noticeable, with EE-MLLM consuming 66G and LLaVA-v1.6 utilizing 69G.
Ablation Study
Implementation Details
Following LLaVA-v1.5 (Liu et al. 2023a), we employ Vicuna-7b-v1.5 (Zheng et al. 2023) as our foundational LLM. The training data is consistent with LLaVA-v1.5, encompassing both pre-training data and supervised fine-tuning data.
Comparisons with LLaVA
In Tab. 4, we compare EE-MLLM with LLaVA under the same settings across both general and fine-grained benchmarks. For the resolution of $336\times 336$, we employ CLIP-ViT-L-14 as the vision encoder, ensuring complete alignment with LLaVA-v1.5. EE-MLLM demonstrates comparable performance to LLaVA on general benchmarks. In terms of fine-grained benchmarks, at $336\times 336$ resolution, EE-MLLM exhibits promising performance on AI2D and ChartQA but slightly underperforms LLaVA on OCRBench and TextVQA. The average score for the $336\times 336$ resolution is 47.1, which is 98% of LLaVA’s 48.1. For the resolution of $672\times 672$, we utilize SigLIP as the vision encoder. As shown in Tab. 4, EE-MLLM obtains comparable results on AI2D and TextVQA, and the average score maintains 98% of LLaVA’s performance.
V O FFN SeedB Hall TextVQA ChartQA DocVQA Avg.
$336\times 336$
65.3 24.4 44.8 14.4 18.4 33.5
66.3 25.9 45.5 16.7 19.2 34.7
65.4 25.3 45.4 15.6 19.1 34.2
65.8 25.1 43.5 14.8 17.9 33.4
66.1 25.7 46.1 15.8 18.6 34.5
$672\times 672$
66.1 25.2 56.8 19.7 28.6 39.3
66.6 25.5 58.2 23.0 29.8 40.6
66.1 24.3 58.2 21.6 30.1 40.1
66.4 24.1 56.3 18.5 26.5 38.4
66.9 27.6 58.8 24.0 31.8 41.8
Table 6: Ablation of the aligner. Checkmarks indicate the weights employed in the aligner.
Ablation of the Aligner
In Tab. 6, we evaluate the aligner variants with different mapping weights. We employ SigLIP initialized from Idefics2 (Laurençon et al. 2024) to conduct experiments at different input resolutions. We remove the different weights from the aligner and train the model at a resolution of $336\times 336$. Checkmarks indicate the weights employed in the aligner. We have three findings: 1) Removing the entire aligner, as depicted in the first row, significantly reduces performance across multiple benchmarks. Specifically, the TextVQA score drops from 46.1 to 44.8, and the average score for the five benchmarks decreases from 34.5 to 33.5. This result highlights the aligner’s effectiveness in aligning visual features with the LLM feature space, allowing text tokens to capture crucial visual information and address questions through causal cross-attention modules. 2) When ablating individual weights in the aligner, we find that maintaining the structure is more critical. The absence of $V$ or $O$ has a relatively minor impact on low-resolution input, even resulting in slightly higher average performance when $V$ is missing. However, when the FFN is missing, the aligner’s structure no longer resembles a transformer block, leading to a significant performance loss. 3) We directly increase the input image resolution to $672\times 672$ without additional training and compare the variants with different aligner types. We observe that, at high-resolution input, the absence of $V$ or $O$ weights results in a substantial decline in fine-grained benchmarks such as TextVQA, ChartQA, and DocVQA. This finding demonstrates the importance of a complete aligner when applied to high-resolution images.
Visualization
We present four examples sampled from BLINK (Fu et al. 2024) and RealWorldQA (XAI 2024) to assess the impact of the architectural changes in Fig 3. The first example demonstrates that EE-MLLM can perceive fine-grained visual context within the image, such as the color of traffic lights. The second and third examples highlight EE-MLLM’s ability to comprehend the position of objects. Specifically, EE-MLLM can accurately identify guitar’s position relative to the table and the dog’s position. The last example reveals that EE-MLLM can distinguish subtle differences in visual content.
Conclusion
In this paper, we revisit previous studies on multimodal large language models and categorize them into two groups: self-attention-based methods and cross-attention-based methods. The former is data-efficient but computationally inefficient, while the latter is compute-efficient but data-inefficient. To ensure data-efficient while maintaining compute-efficient, we propose the composite attention mechanism for EE-MLLM, which incorporates the composite attention module for computational efficiency and the aligner for data efficiency. We conduct comprehensive experiments on general benchmarks and fine-grained benchmarks and find that EE-MLLM achieves state-of-the-art performance on most benchmarks. We also assess the inference speed of EE-MLLM in real-world scenarios, and the results demonstrate that EE-MLLM holds a significant advantage in inference.

### v3

@@ 9 Evalution Benchmarks
We conduct our evaluations using the VLMEvalKit [11], and the results of other state-of-the-art models are obtained from the same source.
9.1 General Benchmarks
@@
1) MMBench-EN [30] is a comprehensive multimodal benchmark specifically designed to assess the performance of MLLMs. It comprises over 3,000 multiple-choice questions spanning 20 ability categories. We evaluate EE-MLLM on the MMBench-EN-V1.1. 2) MME [13] evaluates advanced MLLMs in terms of perception and cognition, encompassing a total of 14 subtasks. To minimize the impact of prompt engineering on MLLMs, MME’s instructions are designed to elicit simple binary responses, such as “please answer yes or no.”. We report the results of the perception split of MME. 3) ScienceQA [34]is derived from elementary and high school science curricula. The questions in ScienceQA cover three subjects: natural science, language science, and social science. 4) HallusionBench [15] is designed for evaluating image-context reasoning, comprising 346 images paired with 1,129 questions crafted by human experts. HallusionBench takes into account both language hallucinations and visual illusions across a diverse range of topics. 5) MMMU [45] collects 11.5K multimodal questions from college exams, quizzes, and textbooks, covering six core disciplines, spanning 30 subjects and 183 subfields, including 30 heterogeneous image types. 6) CCBench [30], developed by the MMBench team, is specifically designed to evaluate MLLMs in the domain of Chinese Culture. 7) SeedBench [23] encompasses 19K multiple choice questions, covering 12 evaluation dimensions, including both image and video. We only use questions with images for evaluation. 8) BLINK [14] contains 14 visual perception tasks which pose significant challenges for current multimodal LLMs.
@@
Figure 6: Inference speed ratios of EE-MLLM and LLaVA for single and double image inputs.
9.2 Fine-grained Benchmarks
1) AI2D [19] emphasizes diagram interpretation and reasoning, comprising 5,000 diagrams and 15,000 questions and answers. 2) OCRBench [31] aims to facilitate the assessment of MLLM OCR capabilities, including 29 datasets. 3) TextVQA [40] consists of 45,336 questions and 28,408 images that necessitate reasoning about text for answering. We use the validation set, containing 5,000 images, for evaluation. 4) ChartQA [36] is a large-scale benchmark featuring 20,882 charts, with questions focusing on logical and visual reasoning. 5) DocVQA [37] concentrates on document image understanding, encompassing 50,000 questions and over 12,000 images. We evaluate using the validation set, which includes 5,349 questions and 1,286 images. 6) Seed2 Plus [24] specifically designed for text-rich visual comprehension evaluation of MLLMs, includes 2.3K multiple-choice questions covering charts, maps, and webs.
10 Comparison of Overall Inference Speed
For MLLMs, the overall inference time consists of the prefilling time and the decode time. The prefilling time mainly involves processing visual inputs and textual instructions, while the decode time refers to the duration required to generate new tokens.
In the main text, we have demonstrated EE-MLLM’s substantial improvements over LLaVA in reducing prefilling time. Here, we further compare the overall inference speed. In Fig. 6, we present the speed ratio between EE-MLLM and LLaVA when generating varying numbers of tokens. We conduct the comparisons on a single NVIDIA H800. The resolution of the input image is set to 980×980, and the number of generated tokens varies from 2 to 256.
@@ 10 Comparison of Overall Inference Speed
Our findings show that when generating 8 tokens, EE-MLLM’s inference speed is three times that of LLaVA. However, as the number of generated tokens increases, the speed ratio decreases. When generating 64 tokens, EE-MLLM’s inference speed is 1.6 times that of LLaVA. The reason for this observation is that our EE-MLLM primarily reduces the computational cost during the prefilling stage, which calculates the KV Cache of visual tokens. The generation of the first token is faster than self-attention-based methods like LLaVA. However, the advantage in inference speed diminishes after the first token. Although the inference speed of EE-MLLM decreases as the number of generated tokens increases, it still maintains an advantage over LLaVA, especially under multi-image input scenarios.